# Trivial Isomorphisms between Reduced Products

### v2

@@ 3. Open Colouring Axioms
In [10], the second author defined an apparent strengthening of $\operatorname{{\mathsf{OCA}_{T}}}$ named $\operatorname{{\mathsf{OCA}_{\infty}}}$. [^14] In [7] a weaker axiom was called $\operatorname{{\mathsf{OCA}_{\infty}}}$, although a proof that $\operatorname{{\mathsf{OCA}_{\infty}}}$ as introduced in [10] is a consequence of $\operatorname{{\mathsf{PFA}}}$ was given. (We should note that this proof contains a glaringly obvious typo; in 4. of Lemma 3.1, ‘whenever $s\subseteq t$’ should be ‘for some $t$ such that $s\subseteq t$.’) This axiom was proved to be equivalent to $\operatorname{{\mathsf{OCA}_{T}}}$ ([30, §5], see also [12, Theorem 8.6.6]). $\operatorname{{\mathsf{OCA}_{\infty}}}$ can be viewed as a refined version of $\operatorname{{\mathsf{OCA}_{\infty}}}$, and Moore’s proof that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}_{\infty}}}$ easily adapts to prove that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$.
@@
Define $Y=\{0,1\}^{\mathbb{N}}\times\prod_{j}{\mathcal{V}}_{j}\times X$. Since each ${\mathcal{V}}_{j}$ is countable, $\prod_{j}{\mathcal{V}}_{j}$ is naturally homeomorphic to the Baire space and $Y$ has a natural Polish space topology. Define a subset $K_{0}$ of $[Y]^{2}$ by letting $\{(a,\mu,x),(b,\nu,y)\}\in K_{0}$ if
@@
Since each element of $\bigcup_{j}{\mathcal{V}}_{j}$ is symmetric, $K_{0}$ is a symmetric subset of $Y^{2}$, clearly disjoint from the diagonal. It is evidently an open subset of $[Y]^{2}$ in its natural Polish topology.
@@ 3.1. A uniformization principle for families of partial selectors
If ${\mathcal{Y}}_{n}$ are topological spaces we topologize the space of partial selectors, and any subset ${\mathcal{F}}$ of it, as follows. Fix a point $0_{n}\in{\mathcal{Y}}_{n}$, and identify $(g,D(g))$ with $\hat{g}\in\prod_{n}{\mathcal{Y}}_{n}$ which agrees with $g$ on $D(g)$ and satisfies $g(n)=0_{n}$ for $n\in{\mathbb{N}}\setminus D(g)$. Via this identification, ${\mathcal{F}}$ is identified with a subset of $\prod_{n}{\mathcal{Y}}_{n}$, which we consider with respect to the (second countable) product topology.
@@
Suppose that $J(s)$, for $s\in\{0,1\}^{<{\mathbb{N}}}$, are disjoint finite subsets of ${\mathbb{N}}$. For $h\in\{0,1\}^{\mathbb{N}}$ let
@@
For $s\in\{0,1\}^{<{\mathbb{N}}}$ let $|s|$ denote its length.
@@
The second part of the following will be applied in situations given by the second alternative of $\operatorname{{\mathsf{OCA}^{\#}}}$.
@@
    Suppose that $S(n,i)$, for $n\in{\mathbb{N}}$ and $i<2^{n}$, are finite sets such that $|S(n,i)|\geq n+(4^{n+1}-1)/3$ for all $n$ and $i$. Then there are pairwise disjoint sets $F(n,i)\subseteq S(n,i)$ such that $|F(n,i)|=2^{n}$ for all $n$ and $i$.
@@
    If $A(s)\in\operatorname{Fin}$ for $s\in\{0,1\}^{<{\mathbb{N}}}$ satisfy $|A(s)|\geq|s|+(4^{n+1}-1)/3$ and $Z\subseteq\{0,1\}^{\mathbb{N}}$ is uncountable without isolated points then there are disjoint $B(s)\subseteq A(s)$ for $s\in\Delta(Z)$ such that $|B(s)|\geq 2^{|s|}$ for all $s$.
@@
(1): If $S(n,i)$ are as described, a fairly dumb algorithm produces the sets $F(n,i)$. Let $F(0,0)$ be any element of $S(0,0)$. Assume that $F(k,i)$ as required had been chosen for all $k<n$ and $i<2^{k}$. Then $I=\bigcup_{k<n,i<2^{k}}F(k,i)\cup n$ has cardinality not greater than $n+\sum_{k<n,i<2^{k}}2^{k}=n+(4^{n}-1)/3$. Therefore $S(n,i)\setminus I$ has cardinality at least $4^{n}$ for each $i<n$. We can now choose disjoint $F(n,i)\subseteq S(n,i)\setminus\bigcup_{j<i}F(n,j)$ of cardinality $2^{n}$ each recursively in $i$.
[…]
@@
Fix $l$. Since ${\mathcal{Y}}_{l}$ is second-countable, let $V(l,m,i)$, $m\in{\mathbb{N}}$, $i=0,1$ of basic open sets such that $V(l,m,0)$ and $V(l,m,1)$ are disjoint and for every pair of distinct points $y_{0},y_{1}$ in ${\mathcal{Y}}_{l}$ there is $m$ such that $y_{i}\in V(l,m,i)$ for $i=0,1$.
@@
Let $W^{n}_{m}=U^{n}_{m,0}\times U^{n}_{m,1}\cup U^{n}_{m,1}\times U^{n}_{m,0}$ and let $\mathcal{W}^{n}=\{W^{n}_{m}\mid m\in{\mathbb{N}}\}$. By $\operatorname{{\mathsf{OCA}_{T}}}$ and Theorem 3.3, $\operatorname{{\mathsf{OCA}^{\#}}}$ holds. Assume that the alternative (1) of $\operatorname{{\mathsf{OCA}^{\#}}}$ applied to $X={\mathcal{F}}$ and the family $\mathcal{W}^{n}$, for $n\in{\mathbb{N}}$, holds. Fix a partition $X=\bigcup_{n}X_{n}$ such that (with $K_{0}^{n}$ as defined in the statement of $\operatorname{{\mathsf{OCA}^{\#}}}$) we have $[X_{n}]^{2}\cap K^{n}_{0}=\emptyset$ for all $n$. Then with ${\mathcal{F}}_{n}=X_{n}$ and all $g,h$ in ${\mathcal{F}}_{n}$ we have $\operatorname{Diff}(g,h)<n+(4^{n+1}-1)/3$, and therefore after re-enumerating the sets ${\mathcal{F}}_{n}$ and interleaving empty sets, the alternative (1) of Proposition 3.6 holds.
@@
$\{f(a_{0}),f(a_{1})\}\in W^{j}_{m(j)}.$
This means (after swapping $a_{0}$ with $a_{1}$ if needed) that $f(a_{i})(l)\in V(j,u^{j}_{m(s)}(l),i)$ for all $l\in S(j,m(s))$ and $i=0,1$. For $s\in\Delta(Z)$ let $A(s)=S(|s|,m(s))$. Then (3.3) implies $\operatorname{Diff}(f(a_{0}),f(a_{1}))\supseteq A(s)$, and $|A(s)|=|s|+(4^{|s|+1}-1)/3$. We may assume that $Z$ has no isolated points by removing all basic open sets $U$ such that $Z\cap U$ is countable from $Z$. Therefore (2) of Lemma 3.7 implies that there are pairwise disjoint sets $B(s)\subseteq A(s)$ such that $|B(s)|\geq 2^{|s|}$, for $s\in\Delta(Z)$. To recap, we have the following.
@@
    There are an uncountable $Z\subseteq\{0,1\}^{\mathbb{N}}$, $f\colon Z\to{\mathcal{F}}$, and pairwise disjoint $B(s)\subset{\mathbb{N}}$ for $s\in\Delta(Z)$ such that $|B(s)|\geq 2^{|s|}$ and for all $x\neq y$ in $Z$ we have $\operatorname{Diff}(f(x),f(y))\supseteq B(x\wedge y)$.
@@
$B_{h}=\{k(s,h\restriction|s|)\mid s\in\{0,1\}^{\mathbb{N}}\}.$
For $h\neq h^{\prime}$ in $\{0,1\}^{\mathbb{N}}$ we have $B_{h}\cap B_{h^{\prime}}\subseteq\bigcup_{|s|<\Delta(h.h^{\prime})}B(s)$, and therefore these sets form a perfect tree-like almost disjoint family ${\mathcal{A}}$. Since for distinct $a$ and $b$ in $Z$ we have $\operatorname{Diff}(f(a),f(b))\supseteq B(a\wedge b)$ and $B(a\wedge b)\cap B_{h}\neq\emptyset$, (2) follows. ∎
@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
The following is Theorem 1 as stated in the introduction and it improves the main result of [39].
@@
We will prove a stronger result along the lines of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem of [8] (Theorem 4.7). The proof of this theorem occupies this section.
@@
Since the constant map sending $\mathcal{P}({\mathbb{N}})$ to the empty set is continuous and lifts $\Phi$ on $\operatorname{Fin}$, ${\mathcal{J}}_{\textrm{cont}}$ contains $\operatorname{Fin}$. We work by contradiction, and assume that ${\mathcal{J}}_{\textrm{cont}}$ is meager. By the well-known characterization of nonmeager ideals of Talagrand and Jalali-Naini ([8, Theorem 3.10.1]), there are disjoint finite subsets, $F_{j}$, for $j\in{\mathbb{N}}$, of ${\mathbb{N}}$ such that $\bigcup_{j\in A}F_{j}\in{\mathcal{J}}_{\textrm{cont}}$ if only if $A$ is finite. Let $F^{\prime}_{i}=\bigcup_{j=2^{i}}^{2^{i+1}-1}F_{j}$. Fix any uncountable almost disjoint family ${\mathcal{A}}_{0}$ in ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ and let
@@
This is a natural tree-like almost disjoint family of cardinality $2^{\aleph_{0}}$ associated to ${\mathcal{A}}$. (Succinctly said, this is the image of ${\mathcal{A}}_{0}$ under the endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ associated with the Rudin–Blass reduction of $\operatorname{Fin}$ to itself corresponding to the function that collapses $F_{i}^{\prime}$ to $i$ for all $i\in{\mathbb{N}}$.)
The next step of the proof is virtually identical to that of [39, Lemma 2.2], where it was proved that $\operatorname{{\mathsf{OCA}_{T}}}$ implies that all but countably many of the sets in ${\mathcal{A}}$ belong to ${\mathcal{J}}_{\sigma}$. Note, in [39, Lemma 2.2] the assumption that $\Phi$—denoted $\phi$ in [39]—is an automorphism isn’t used until after $\Phi$ is proven to have a $\sigma$-Borel-measurable lifting on ${\mathcal{P}}(B)$ for some set $B$ in the given tree-like almost disjoint family, at which point [39, Theorem 1.2] is invoked. As this theorem relies on the assumption that $\Phi$ is an automorphism, at this point we depart from the proof of [39, Lemma 2.2] and instead proceed as follows.
A set $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\sigma}$ is of the form $\bigcup_{n\in A}F_{n}’=\bigcup_{n\in A}\bigcup_{j=2^{n}}^{2^{n+1}}F_{j}$ for some $A\in{\mathcal{A}}_{0}$. Then $B_{i}=\bigcup_{j\in A,j<2^{i}}F_{2^{i}+j}$, for $i\in{\mathbb{N}}$, is a partition of $B$ into sets each one of which is an infinite union of the original intervals $F_{j}$, and it therefore does not belong to ${\mathcal{J}}_{\textrm{cont}}$. However, Proposition A.2 proven in the Appendix implies that all but finitely many of the $B_{i}$ belong to ${\mathcal{J}}_{\textrm{cont}}$; contradiction. ∎
@@
It is straightforward to verify that $\{A,B\}\in L_{1}(D)$ if and only if $h_{A}$ and $h_{B}$ agree on $h_{A}^{-1}(D)\cap h_{B}^{-1}(D)$. By identifying $A$ with the graph of $g_{A}$, we see that $\operatorname{{\mathsf{OCA}_{T}}}$ applies to this partition.
@@
${\mathcal{X}}_{n}=\{A\in{\mathcal{J}}_{\textrm{cont}}|g_{A}(m)=g_{D}(m)\text{ for all }m\in(A\cap D)\setminus(n+1)\}.$
@@
A function between Polish spaces is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets [^16] Kechris in [26, §29.D] calls this functions $\sigma(\Sigma^{1}_{1})$-measurable, while C-measurability denotes a weaker condition. Nevertheless, we stick to this terminology as it is commonly used in the area of study of liftings (e.g., [12, §17], [29], or [11]).. In the following proof, we use the Jankov–von Neumann uniformization theorem, stated for convenience in Appendix A.
@@
For every $A\in{\mathcal{J}}_{\textrm{cont}}$, fix $g_{A}\colon A\to\operatorname{Fin}$ such that $B\mapsto\bigcup_{n\in B}g_{A}(n)$ is a lifting of $\Phi$ on ${\mathcal{P}}(A)$. Let
=${\mathcal{F}}=\{(A,g_{A})\mid A\in{\mathcal{J}}_{\textrm{cont}}\}.$
This is a family of partial selectors, with ${\mathcal{Y}}_{n}=\operatorname{Fin}^{\mathbb{N}}$ for all $n$ (see §3.1; this proof makes heavy use of terminology and results from this section). $\operatorname{{\mathsf{OCA}^{\#}}}$ and Proposition 3.6 together give us two alternatives.
If (2) of Proposition 3.6 applies, then there are a perfect tree-like almost disjoint family ${\mathcal{A}}$, an uncountable $Z\subseteq\{0,1\}^{\mathbb{N}}$, and a continuous $f\colon Z\to{\mathcal{F}}$ such that for every $A\in{\mathcal{A}}$ and all distinct $x,y$ in $Z$ we have
@@
By Proposition 4.3 and $\operatorname{{\mathsf{OCA}_{T}}}$, ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager. It therefore intersects ${\mathcal{A}}$ nontrivially, by the same result of Talagrand and Jalali-Naini used in the proof of Proposition 4.3 ([8, Theorem 3.10.1]). If $A\in{\mathcal{A}}\cap{\mathcal{J}}_{\textrm{cont}}$ then Lemma 4.6 implies that ${\mathcal{J}}_{\textrm{cont}}$ is $\sigma$-$L_{1}(A)$-homogeneous. However, $f[Z]$ is an uncountable $L_{0}(A)$-homogeneous subset of ${\mathcal{J}}_{\textrm{cont}}$; contradiction.

### v3

@@ 3. Open Colouring Axioms
In [10], the second author defined an apparent strengthening of $\operatorname{{\mathsf{OCA}_{T}}}$ named $\operatorname{{\mathsf{OCA}_{\infty}}}$. [^14] In [7] a weaker axiom was called $\operatorname{{\mathsf{OCA}_{\infty}}}$, although a proof that $\operatorname{{\mathsf{OCA}_{\infty}}}$ as introduced in [10] is a consequence of $\operatorname{{\mathsf{PFA}}}$ was given. (We should note that this proof contains a glaringly obvious typo; in 4. of Lemma 3.1, ‘whenever $s\subseteq t$’ should be ‘for some $t$ such that $s\subseteq t$.’) This axiom was proved to be equivalent to $\operatorname{{\mathsf{OCA}_{T}}}$ ([30, §5], see also [12, Theorem 8.6.6]). $\operatorname{{\mathsf{OCA}^{\#}}}$ can be viewed as a refined version of $\operatorname{{\mathsf{OCA}_{\infty}}}$, and Moore’s proof that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}_{\infty}}}$ easily adapts to prove that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$.
@@
Define $Y=\{0,1\}^{\mathbb{N}}\times\prod_{j}{\mathcal{V}}_{j}\times X$. Since each ${\mathcal{V}}_{j}$ is countable, $\prod_{j}{\mathcal{V}}_{j}$ is naturally homeomorphic to the Baire space and $Y$ has a natural separable metrizable topology. Define a subset $K_{0}$ of $[Y]^{2}$ by letting $\{(a,\mu,x),(b,\nu,y)\}\in K_{0}$ if
@@
Since each element of $\bigcup_{j}{\mathcal{V}}_{j}$ is symmetric, $K_{0}$ is a symmetric subset of $Y^{2}$, clearly disjoint from the diagonal. It is evidently an open subset of $[Y]^{2}$ in its natural topology.
@@ 3.1. A uniformization principle for families of partial selectors
If ${\mathcal{Y}}_{n}$ are topological spaces we then topologize the space of partial selectors as follows. Fix a point $0_{n}\in{\mathcal{Y}}_{n}$, and identify $(g,D(g))$ with $\hat{g}\in\prod_{n}{\mathcal{Y}}_{n}$ which agrees with $g$ on $D(g)$ and satisfies $g(n)=0_{n}$ for $n\in{\mathbb{N}}\setminus D(g)$. Via this identification, a set ${\mathcal{F}}$ of partial selectors is identified with a subset of $\prod_{n}{\mathcal{Y}}_{n}$ and equipped with the subspace topology.
In our applications all ${\mathcal{Y}}_{n}$ will be separable and metrizable (second countable), hence the topology on ${\mathcal{F}}$ will also be separable and metrizable (second countable). If all spaces ${\mathcal{Y}}_{n}$ are Polish then so is the space of partial selectors.
@@
Suppose that $J(s)$, for $s\in\{0,1\}^{<{\mathbb{N}}}$, are pairwise disjoint nonempty finite subsets of ${\mathbb{N}}$. For $h\in\{0,1\}^{\mathbb{N}}$ let
@@
For $s\in\{0,1\}^{<{\mathbb{N}}}$ let $|s|$ denote its length (equivalently, its domain).
@@
The second part of the following will be applied in situations given by the second alternative of $\operatorname{{\mathsf{OCA}^{\#}}}$, with $D=\Delta(Z)$ for a carefully chosen uncountable subset $Z$ of the Cantor set.
@@
    Suppose that $S(n,i)$, for $n\in{\mathbb{N}}$ and $i<2^{n}$, are finite sets such that $|S(n,i)|\geq n+(4^{n+1}-1)/3$ for all $n$ and $i$. Then there are pairwise disjoint sets $F(n,i)\subseteq S(n,i)\setminus n$ such that $|F(n,i)|=2^{n}$ for all $n$ and $i$.
@@
    If $A(s)\in\operatorname{Fin}$ for $s\in D\subseteq\{0,1\}^{<{\mathbb{N}}}$ satisfy $|A(s)|\geq|s|+(4^{|s|+1}-1)/3$ for all $s\in D$, then there are pairwise disjoint $B(s)\subseteq A(s)\setminus|s|$ for $s\in D$ such that $|B(s)|=2^{|s|}$ for all $s\in D$.
@@
(1): If $S(n,i)$ are as described, a fairly dumb algorithm produces the sets $F(n,i)$. Let $F(0,0)$ be any element of $S(0,0)$. Assume that $F(k,i)$ as required had been chosen for all $k<n$ and $i<2^{k}$. Then $I=\bigcup_{k<n,i<2^{k}}F(k,i)\cup n$ has cardinality not greater than $n+\sum_{k<n,i<2^{k}}2^{k}=n+(4^{n}-1)/3$. Therefore $S(n,i)\setminus I$ has cardinality at least $4^{n}$ for each $i<n$. We can now choose pairwise disjoint $F(n,i)\subseteq S(n,i)\setminus I$ of cardinality $2^{n}$ each recursively in $i$.
(2): Enumerate the $n$th level of $\{0,1\}^{<{\mathbb{N}}}$ as $s(n,i)$, for $i<2^{n}$. Define the sets $S(n,i)$, for $n\in{\mathbb{N}}$ and $i<2^{n}$ by setting $S(n,i)=A(s(n,i))$ when $s(n,i)\in D$ and by choosing $S(n,i)$ to be an arbitrary set of cardinality $n+(4^{n+1}-1)/3$ when $s(n,i)\notin D$. By part (1) of this lemma, there are pairwise disjoint $F(n,i)\subseteq S(n,i)\setminus n$ for $n\in{\mathbb{N}}$ and $i<2^{n}$, of cardinality $2^{n}$. Then by defining for every $s\in D$ the set $B(s)=F(n,i)$, where $n$ and $i<2^{n}$ are such that $s=s(n,i)$, we find that the sets $B(s)\subseteq A(s)\setminus|s|$ are as required.
@@
Fix $l$. Since ${\mathcal{Y}}_{l}$ is second-countable, let $V(l,m,i)$, for $m\in{\mathbb{N}}$ and $i=0,1$, be basic open sets such that $V(l,m,0)$ and $V(l,m,1)$ are disjoint and for every pair of distinct points $y_{0},y_{1}$ in ${\mathcal{Y}}_{l}$ there is $m$ such that $y_{i}\in V(l,m,i)$ for $i=0,1$.
@@
Let $W^{n}_{m}=U^{n}_{m,0}\times U^{n}_{m,1}\cup U^{n}_{m,1}\times U^{n}_{m,0}$ and let $\mathcal{W}^{n}=\{W^{n}_{m}\mid m\in{\mathbb{N}}\}$. By $\operatorname{{\mathsf{OCA}_{T}}}$ and Theorem 3.3, $\operatorname{{\mathsf{OCA}^{\#}}}$ holds. Assume that the alternative (1) of $\operatorname{{\mathsf{OCA}^{\#}}}$ applied to $X={\mathcal{F}}$ and the family $\mathcal{W}^{n}$, for $n\in{\mathbb{N}}$, holds. Fix a partition $X=\bigcup_{n}X_{n}$ such that we have $[X_{n}]^{2}\cap\bigcup\mathcal{W}^{n}=\emptyset$ for all $n$. Then with ${\mathcal{F}}_{n}=X_{n}$, for all $g,h$ in ${\mathcal{F}}_{n}$ we have $\operatorname{Diff}(g,h)<n+(4^{n+1}-1)/3$, and therefore after re-enumerating the sets ${\mathcal{F}}_{n}$ and interleaving empty sets, the alternative (1) of Proposition 3.7 holds.
@@
$\{f(a_{0}),f(a_{1})\}\in W^{j}_{m(s)}.$
This means (after swapping $a_{0}$ with $a_{1}$ if needed) that $f(a_{i})(l)\in V(j,u^{j}_{m(s)}(l),i)$ for all $l\in S(j,m(s))$ and $i=0,1$. For $s\in\Delta(Z)$ let $A(s)=S(|s|,m(s))$. Then (3.3) implies $\operatorname{Diff}(f(a_{0}),f(a_{1}))\supseteq A(s)$, and $|A(s)|=|s|+(4^{|s|+1}-1)/3$. We may assume that $Z$ has no isolated points by removing all basic open sets $U$ such that $Z\cap U$ is countable from $Z$. Furthermore, (2) of Lemma 3.8 implies that there are pairwise disjoint sets $B(s)\subseteq A(s)\setminus|s|$ such that $|B(s)|=2^{|s|}$, for $s\in\Delta(Z)$. To recap, we have the following.
@@
    There are an uncountable $Z\subseteq\{0,1\}^{\mathbb{N}}$, $f\colon Z\to{\mathcal{F}}$, and pairwise disjoint $B(s)\subset{\mathbb{N}}$ for $s\in\Delta(Z)$ such that $|B(s)|=2^{|s|}$ and for all $x\neq y$ in $Z$ we have $\operatorname{Diff}(f(x),f(y))\setminus\Delta(x,y)\supseteq B(x\wedge y)$.
@@
$B_{h}=\{k(s,h\restriction|s|)\mid s\in\Delta(Z)\}.$
For $h\neq h^{\prime}$ in $\{0,1\}^{\mathbb{N}}$ we have $B_{h}\cap B_{h^{\prime}}\subseteq\bigcup_{|s|<\Delta(h.h^{\prime})}B(s)$, and therefore these sets form a perfect tree-like almost disjoint family ${\mathcal{A}}$. Since for distinct $x$ and $y$ in $Z$ we have $\operatorname{Diff}(f(x),f(y))\setminus\Delta(x,y)\supseteq B(x\wedge y)$ and $B(x\wedge y)\cap B_{h}\neq\emptyset$, (2) follows. ∎
@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
The following is the first part of Theorem 1 as stated in the introduction.
@@
We will prove a stronger result along the lines of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem of [8] (see Theorem 4.7, which is the second part of Theorem 1 as stated in the introduction). The proof of this theorem occupies this section.
@@
Since the constant map sending $\mathcal{P}({\mathbb{N}})$ to the empty set is continuous and lifts $\Phi$ on $\operatorname{Fin}$, ${\mathcal{J}}_{\textrm{cont}}$ contains $\operatorname{Fin}$. We work by contradiction, and assume that ${\mathcal{J}}_{\textrm{cont}}$ is meager. By the well-known characterization of nonmeager ideals of Talagrand and Jalali-Naini ([8, Theorem 3.10.1]), there are pairwise disjoint finite intervals of ${\mathbb{N}}$, $F_{j}$, for $j\in{\mathbb{N}}$, such that $\bigcup_{j\in A}F_{j}\in{\mathcal{J}}_{\textrm{cont}}$ if and only if $A$ is finite. Let $F^{\prime}_{i}=\bigcup_{j=2^{i}}^{2^{i+1}-1}F_{j}$. Fix any perfect tree-like almost disjoint family ${\mathcal{A}}_{0}$ and let
@@
This is a natural perfect tree-like almost disjoint family associated to ${\mathcal{A}}_{0}$. (Succinctly said, this is the image of ${\mathcal{A}}_{0}$ under the endomorphism of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ associated with the Rudin–Blass reduction of $\operatorname{Fin}$ to itself corresponding to the function that collapses $F_{i}^{\prime}$ to $i$ for all $i\in{\mathbb{N}}$.)
The next step of the proof is virtually identical to that of [38, Lemma 2.2], where it was proved that $\operatorname{{\mathsf{OCA}_{T}}}$ implies that all but countably many of the sets in ${\mathcal{A}}$ belong to ${\mathcal{J}}_{\sigma}$. Note, in [38, Lemma 2.2] the assumption that $\Phi$—denoted $\varphi$ in [38]—is an automorphism isn’t used until after $\Phi$ is proven to have a $\sigma$-Borel-measurable lifting on ${\mathcal{P}}(B)$ for some set $B$ in the given tree-like almost disjoint family, at which point [38, Theorem 1.2] is invoked. As this theorem relies on the assumption that $\Phi$ is an automorphism, at this point we depart from the proof of [38, Lemma 2.2] and instead proceed as follows.
A set $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\sigma}$ is of the form $\bigcup_{i\in A}F_{i}’=\bigcup_{i\in A}\bigcup_{j=2^{i}}^{2^{i+1}-1}F_{j}$ for some $A\in{\mathcal{A}}_{0}$. Then $B_{j}=\bigcup_{i\in A,j<2^{i}}F_{2^{i}+j}$, for $j\in{\mathbb{N}}$, is a partition of $B$ into sets each one of which is an infinite union of the original intervals $F_{j}$, and it therefore does not belong to ${\mathcal{J}}_{\textrm{cont}}$. However, Proposition A.2 proven in the Appendix implies that all but finitely many of the $B_{j}$ belong to ${\mathcal{J}}_{\textrm{cont}}$; contradiction. ∎
@@
By identifying $A$ with the graph of $g_{A}$, we see that $\operatorname{{\mathsf{OCA}_{T}}}$ applies to the partition $[{\mathcal{J}}_{\textrm{cont}}]^{2}=L_{0}(D)\cup L_{1}(D)$.
@@
${\mathcal{X}}_{n}=\{A\in{\mathcal{J}}_{\textrm{cont}}\mid g_{A}(m)=g_{D}(m)\text{ for all }m\in(A\cap D)\setminus(n+1)\}.$
@@
A function between Polish spaces is called C-measurable if it is measurable with respect to the $\sigma$-algebra generated by analytic sets [^16] Kechris in [26, §29.D] calls such functions $\sigma(\Sigma^{1}_{1})$-measurable, while C-measurability denotes a weaker condition. Nevertheless, we stick to this terminology as it is commonly used in the area of study of liftings (e.g., [12, §17], [29], or [11]).. In the following proof, we use the Jankov–von Neumann uniformization theorem, stated for convenience in Appendix A.
@@
Recall that for every $A\in{\mathcal{J}}_{\textrm{cont}}$ we have fixed $g_{A}\colon A\to\operatorname{Fin}$ such that $B\mapsto\bigcup_{n\in B}g_{A}(n)$ is a lifting of $\Phi$ on ${\mathcal{P}}(A)$. Let
=${\mathcal{F}}=\{(g_{A},A)\mid A\in{\mathcal{J}}_{\textrm{cont}}\}.$
This is a family of partial selectors, with ${\mathcal{Y}}_{n}=\operatorname{Fin}$ for all $n$ (see §3.1; this proof makes heavy use of terminology and results from this section). $\operatorname{{\mathsf{OCA}_{T}}}$ and Proposition 3.7 together give us two alternatives.
If (2) of Proposition 3.7 applies, then there are a perfect tree-like almost disjoint family ${\mathcal{A}}$, an uncountable $Z\subseteq\{0,1\}^{\mathbb{N}}$, and a function $f\colon Z\to{\mathcal{F}}$ such that for every $A\in{\mathcal{A}}$ and all distinct $x,y$ in $Z$ we have
@@
By Proposition 4.3 and $\operatorname{{\mathsf{OCA}_{T}}}$, ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager. It therefore intersects ${\mathcal{A}}$ nontrivially, by the same result of Talagrand and Jalali-Naini used in the proof of Proposition 4.3 ([8, Theorem 3.10.1]). If $A\in{\mathcal{A}}\cap{\mathcal{J}}_{\textrm{cont}}$ then Lemma 4.6 implies that ${\mathcal{J}}_{\textrm{cont}}$ is $\sigma$-$L_{1}(A)$-homogeneous. However, if $x$ and $y$ are distinct elements of $Z$ with $f(x)=(g_{A_{1}},A_{1})$ and $f(y)=(g_{A_{2}},A_{2})$, then the existence of an element of $\operatorname{Diff}(g_{A_{1}},g_{A_{2}})\cap A$ gives that $\{A_{1},A_{2}\}\in L_{0}(A)$. This would give that $f[Z]$ is an uncountable $L_{0}(A)$-homogeneous subset of the $\sigma$-$L_{1}(A)$-homogeneous set ${\mathcal{J}}_{\textrm{cont}}$, which is a contradiction.